OpenStudy (faman39):

A wedding cake is prepared using 3 individual cakes. The first is a rectangular prism with a length 11 inches, a width 7 inches, and a height 2 inches. The second is a cylinder with a diameter 8 inches and a height 3 inches. The third is a cube with side length 5 inches. What is the approximate total volume of the cake? Use 3.14 for pi . Round to the nearest whole number. 330 in3 430 in3 471 in3 882 in3

OpenStudy (lukebluefive):

Volume of rectangular prism = length x width x height = 11 inches x 7 inches x 2 inches Volume of cylinder = height x PI x radius^2 = 3 inches x 3.14 x (8/2)^2 inches Volume of cube = side^3 = 5^3 inches

OpenStudy (lukebluefive):

Sum all three together and you'll have your approximate answer. Remember to round the cylinder's volume to the nearest integer before adding it to the others.

OpenStudy (faman39):

Thank you soooooooooooooooo much @LukeBlueFive :)

OpenStudy (lukebluefive):

@Faman39 You're very welcome!

OpenStudy (anonymous):

so whats the answer don't tell me how just tell me the answer